OpenStudy (precal):

limit as x approaches 3 from the right (x-3)^(1/2) over (x-1)^(1/2)

OpenStudy (precal):

\[\frac{ \sqrt{x-3} }{ \sqrt{x-1} }\] I need the limit of this function as x approaches 3 from the right

OpenStudy (anonymous):